# A Uniqueness Theorem for Distributed Computation under Physical Constraints

### v1

@@ 6.1. Convergence
Proof Sketch.
The theoretical guarantee for this deterministic outcome in a chaotic environment comes from the semilattice algebraic structure of the systems state space (Axiom 4). The core idea is that every merge operation ($\oplus$) causes the system state to monotonically move upward in the semilattices partial order. The commutativity and associativity of $\oplus$ guarantee independence from order, while idempotency guarantees independence from repetition. Since all execution paths advance monotonically towards a shared upper bound, they must converge to the same unique limit. The detailed formal proof is provided in Appendix F. ∎
6.2. Completeness
@@
Proof Sketch.
This lemma is the cornerstone for proving Turing completeness. It shows that SDPFs core primitives (Branch, Copy, Combine) are sufficient to simulate an atomic find-and-replace step of the SK calculus. The Branch primitive finds a reducible pattern, while Combine and Copy replace it by atomically injecting new tiles and retiring old ones. The robustness of this simulation against network chaos is guaranteed by the statelessness and idempotency of SDPF operations. The detailed construction is provided in Appendix G. ∎
Figure 3. Simulating an SK-combinator reduction ($Sxyz\to xz(yz)$) using SDPF primitives. The transformation is guaranteed to be atomic and idempotent by the Combine operation and its unique rid.
@@
Proof Sketch.
This proof elevates the one-step simulation of Lemma 6.4 to the simulation of any full computation. Lemma 6.4 guarantees the ability to execute each step correctly, while the fairness assumption in Axiom 2 guarantees that any ready step will eventually be executed. This combination ensures that SDPF can simulate any SK reduction sequence. Since SK calculus is Turing-complete, SDPF is as well. The detailed formal proof is provided in Appendix H. ∎
6.3. Minimality
@@
The absence of any metadata field in (id, target, op, pc/next, rid) will either break the feasibility under the trilemma or violate the properties of Theorem 6.2 or 6.5.
Proof Sketch.
This proof establishes that the five metadata fields are the minimal set required to uphold the paradigms guarantees. The strategy is a case-by-case proof by contradiction. We show that removing any single field would fatally undermine a previously established theoretical pillar: absence of id or pc/next breaks completeness; absence of target or op breaks correctness and convergence; and absence of rid breaks idempotency, forcing the system into a high-cost mode that violates the trilemma. Since each field is indispensable, the set is minimal. The detailed formal proof is provided in Appendix I. ∎
7. Discussion
@@
7.1. A Constructive Duality to Impossibility Theorems
@@
7.2. SDPF as a Foundational Model of Computation
@@
7.3. Manifestation in Physically-Realizable Systems
The theoretical framework developed in this paper finds its direct manifestation in the design of physically-realizable systems. The principles of SDPF do not merely align well with programmable data planes; rather, they represent a formalization of the necessary design patterns that emerge under the physical constraints of environments like INC. The data-centric model, with self-describing packets and stateless processing functions, is a direct reflection of the architectural paradigms found in systems like P4 (Bosshart et al., 2014). Our work provides the formal justification for why these architectural choices are optimal and, in fact, necessary.(Fang et al., 2023; Liu et al., 2023; Khashab et al., 2024)
8. Conclusion
@@
Acknowledgements.
@@
Conclusion: Any violation of S2 or S3 forces a choice that violates either Axiom 1 or the linear scalability goal. Thus, S2 and S3 are necessary. ∎
@@ Conclusion
The final paradigm $\mathcal{E}_{4}=T(\mathcal{E}_{0})$ satisfies properties S1-S5 by construction and is therefore in the SDPF normal form. Since we have shown that $\mathcal{E}_{0}\approx\mathcal{E}_{1}\approx\mathcal{E}_{2}\approx\mathcal{E}_{3}\approx\mathcal{E}_{4}$, by transitivity, any optimal paradigm $\mathcal{E}_{0}$ is metric-equivalent to the SDPF normal form. This proves the uniqueness of SDPF. ∎
@@
  3. (3)
    Part 2: Existence of an Upper Bound: Any state $A_{k}$ is the join of a subset of increments from $\Delta^{\star}$. By semilattice properties, the join of a subset is $\preceq$ the join of the superset, so $A_{k}\preceq A_{\infty}$. Thus, the monotonic sequence is bounded above and must converge to a limit. (Tarski, 1955)
  4. (4)
@@
Conclusion: As we have constructed a finite and robust SDPF simulation for the basis combinators, any single-step SK reduction can be simulated. ∎
@@
Conclusion: Since the absence of any single field leads to the failure of at least one core property, the set is minimal. ∎

### v2

@@ 6.1 Convergence
Proof 6.3 (Proof Sketch).
The theoretical guarantee for this deterministic outcome in a chaotic environment comes from the semilattice algebraic structure of the systems state space (Axiom 4). The core idea is that every merge operation ($\oplus$) causes the system state to monotonically move upward in the semilattices partial order. The commutativity and associativity of $\oplus$ guarantee independence from order, while idempotency guarantees independence from repetition. Since all execution paths advance monotonically towards a shared upper bound, they must converge to the same unique limit. The detailed formal proof is provided in F.
6.2 Completeness
@@
Proof 6.6 (Proof Sketch).
This lemma is the cornerstone for proving Turing completeness. It shows that SDPFs core primitives (Branch, Copy, Combine) are sufficient to simulate an atomic find-and-replace step of the SK calculus. The Branch primitive finds a reducible pattern, while Combine and Copy replace it by atomically injecting new tiles and retiring old ones. The robustness of this simulation against network chaos is guaranteed by the statelessness and idempotency of SDPF operations. The detailed construction is provided in G.
Figure 3: Simulating an SK-combinator reduction ($Sxyz\to xz(yz)$) using SDPF primitives. The transformation is guaranteed to be atomic and idempotent by the Combine operation and its unique rid.
@@
Proof 6.8 (Proof Sketch).
This proof elevates the one-step simulation of Lemma 6.5 to the simulation of any full computation. Lemma 6.5 guarantees the ability to execute each step correctly, while the fairness assumption in Axiom 2 guarantees that any ready step will eventually be executed. This combination ensures that SDPF can simulate any SK reduction sequence. Since SK calculus is Turing-complete, SDPF is as well. The detailed formal proof is provided in H.
6.3 Minimality
@@
The absence of any metadata field in (id, target, op, pc/next, rid) will either break the feasibility under the trilemma or violate the properties of Theorem 6.2 or 6.7.
Proof 6.10 (Proof Sketch).
This proof establishes that the five metadata fields are the minimal set required to uphold the paradigms guarantees. The strategy is a case-by-case proof by contradiction. We show that removing any single field would fatally undermine a previously established theoretical pillar: absence of id or pc/next breaks completeness; absence of target or op breaks correctness and convergence; and absence of rid breaks idempotency, forcing the system into a high-cost mode that violates the trilemma. Since each field is indispensable, the set is minimal. The detailed formal proof is provided in I.
7 Discussion
@@
7.1 A Constructive Duality to Impossibility Theorems
@@
7.2 SDPF as a Foundational Model of Computation
@@
7.3 Manifestation in Physically-Realizable Systems
The theoretical framework developed in this paper finds its direct manifestation in the design of physically-realizable systems. The principles of SDPF do not merely align well with programmable data planes; rather, they represent a formalization of the necessary design patterns that emerge under the physical constraints of environments like INC. The data-centric model, with self-describing packets and stateless processing functions, is a direct reflection of the architectural paradigms found in systems like P4 [3]. Our work provides the formal justification for why these architectural choices are optimal and, in fact, necessary.[21, 22, 23]
8 Conclusion
@@
In this work, we focus on deterministic guarantees for the SDPF normal form. The exploration of probabilistic data structures (e.g., Bloom filters) to further trade-off strict correctness for memory efficiency in extreme scale is a promising direction for future research.
[…]
@@
Conclusion: Any violation of S2 or S3 forces a choice that violates either Axiom 1 or the linear scalability goal. Thus, S2 and S3 are necessary.
@@ Conclusion
The final paradigm $\mathcal{E}_{4}=T(\mathcal{E}_{0})$ satisfies properties S1-S5 by construction and is therefore in the SDPF normal form. Since we have shown that $\mathcal{E}_{0}\approx\mathcal{E}_{1}\approx\mathcal{E}_{2}\approx\mathcal{E}_{3}\approx\mathcal{E}_{4}$, by transitivity, any optimal paradigm $\mathcal{E}_{0}$ is metric-equivalent to the SDPF normal form. This proves the uniqueness of SDPF.
@@
  3. 3.
    Part 2: Existence of an Upper Bound: Any state $A_{k}$ is the join of a subset of increments from $\Delta^{\star}$. By semilattice properties, the join of a subset is $\preceq$ the join of the superset, so $A_{k}\preceq A_{\infty}$. Thus, the monotonic sequence is bounded above and must converge to a limit. [24]
  4. 4.
@@
Conclusion: As we have constructed a finite and robust SDPF simulation for the basis combinators, any single-step SK reduction can be simulated.
@@
Conclusion: Since the absence of any single field leads to the failure of at least one core property, the set is minimal.